Prove that SL(n,F) is a normal subgroup of GL(n,F)?


1
Expert's answer
2021-06-29T10:35:09-0400

Recall:

GL(n,R) is the collection of the general

n×n invertible matrices.

SL(n,R)={X∈GL(n,R)∣det(X)=1}.

To prove: SL(n,R) is a normal subgroup of G

Let X∈SL(n,R)

and let P∈G

Then we have

det(PXP-1)=det(P)det(X)det(P)-1=det(X)=1

and hence the conjugate PXP-1

PXP-1 is in SL(n,R)

Therefore, SL(n,R)

SL(n,R) is a normal subgroup of G